C=100+1/2 y^d





I=50-25r





G=100=T





Find equilibrium output assuming interest rate is 15%.





Find IS curve equation and show graphically





Find slope of IS curve.





Government has decided to use expansionary fiscal policy by increasing government spending and tax to 150?




1
Expert's answer
2022-04-04T09:32:14-0400

"C= 100+\frac{1}{2}y^d"

"I= 5-25r"

"G=100 = T"

"G= 100"

"T= 100"

"Y= C+I+G"

"100+0.5YD+ 50-25r+100"

"250+0.5YD-25 r"

"250+0.5(Y-100)-25r"

"250+0.5Y-50-25r"

0"0.5Y = 200-25r"

"Y= 400-50r"

"Y= 392.5"


"250+0.5(Y-150)-25r"

"250+0.5Y-75-25r"

"175+0.5Y-75-25r"

"0.5+0.75-25r"

"Y= 350-50r"

"Y= 350-50r"

"Y= 350-(0.15) 50"

"Y= 342.5"